OpenStudy (anonymous):

Solve: x2 + 6x + 8 = 0 a) x = -2 b) x = -4, -2 c) x = 4, 2 d) x = 4

OpenStudy (anonymous):

Sorry, sorry. B. :D x^2 + 6x + 8 = 0 (x + 4)(x + 2) = 0 x + 4 = 0 x = -4 x + 2 = 0 x = -2